Determine the rule for the sequence:6;-4;6;1;6;6;6;11;6
1
Expert's answer
2020-02-10T09:46:22-0500

Let the sequence be a1, a2, a3, a4, a5, a6, a7, a8, a9, n = "(\overline{1,9})"

If n is odd, an = 6.

Otherwise, an = -4 + 5*("\frac{n}{2}" -1)

According to these rules:

a1 = 6

a2 = -4 + 5*(1-1) = -4

a3 = 6

a4 = -4 + 5*(2-1) = 1

a5 = 6

a6 = -4 + 5*(3-1) = 6

a7 = 6

a8 = -4 + 5*(4-1) = 11

a9 = 6